# Enhancing Redundancy-based Automated Program Repair by Fine-grained Pattern Mining

### v1

@@ 6. Discussion
Threats: The threats to external validity mainly lie in the subjects used in our study. Since the Defects4J benchmark has been widely used by existing studies and contains more than 800 real-world bugs from 17 diverse projects, we believe the results can be reliable. However, the effectiveness of our approach on a broader range remains to be investigated. The threats to internal validity mainly lie in the implementation of our approach. To ensure the correctness and reliability of the experimental results, two authors have carefully checked the correctness of the implementation via code review. In addition, we also publish all our experimental results for replication and promoting future research.
7. Related Work
@@
There are many automated program approaches have been proposed and achieved good results (Weimer et al., 2009; Long et al., 2017; Weimer et al., 2013; Pei et al., 2014; B. Le et al., 2016; Mechtaev et al., 2015; Nguyen et al., 2013; Mechtaev et al., 2018; Jiang et al., 2018; Xiong et al., 2017; Xuan et al., 2017; Le et al., 2016; Ghanbari and Zhang, 2018; Jiang et al., 2023a, 2021), among which redundancy-based APR approaches have been attracted much attention, such as GenProg (Weimer et al., 2009; Le Goues et al., 2012), AE (Weimer et al., 2013; Yuan and Banzhaf, 2018), RSRepair (Qi et al., 2014), ARJA (Yuan and Banzhaf, 2018), ssFix (Xin and Reiss, 2017b), SimFix (Jiang et al., 2018), CapGen (Wen et al., 2018), SearchRpepair (Ke et al., 2015), HERCULES (Saha et al., 2019) and so on. All these approaches are based on the “plastic surgery hypothesis” (Barr et al., 2014). The major difference among them lies in the adopted patch generation strategies. For example, GenProg reuses existing code randomly, ssFix searches similar code as references from a large codebase, while SimFix further incorporates history patches to refine the patch space. Similarly, GapGen considers the frequency of tokens for patch generation while HERCULES leverages the co-evolution of program elements. As explained in the introduction, our approach aims to improve the performance of redundancy-based APR by exploring finer-grained code reuse, which is different from existing approaches. Specifically, compared with SimFix and ssFix, our approach leverages finer-grained token-level code patterns while SimFix and ssFix cannot as they depend on the AST difference between buggy and referenced code. Similarly, our approach is also different from CapGen as Repatt generates patches by mining code patterns while CapGen only considers the frequency of tokens without considering their combinations. In addition, compared with SearchRepair, our approach does not rely on the heavy semantic code search process. In particular, our approach can leverage nonconsecutive token usage patterns for patch generation, which has been proved to be effective but is not supported by these existing approaches. Additionally, PAR (Kim et al., 2013) and TBar (Liu et al., 2019) aim to improve the patch quality via pre-defining a set of repair templates. Apart from those, a set of APR techniques leverage the advance of mature constraint-solving techniques for patch generation (Nguyen et al., 2013; Mechtaev et al., 2015, 2016, 2018; Gao et al., 2021; Xuan et al., 2017; Chen et al., 2017; Hua et al., 2018). However, since these approaches depend on symbolic execution and constraint solving, they usually face the scalability issue for practical use. Different from these approaches, this work aims to improve the performance of redundancy-based APR techniques and does not face the scalability issue.
@@
With the rapid development of deep learning techniques, the latest APR approaches also leverage such techniques for patch generation. Early studies use statistical machine learning algorithms to sort or pick patch ingredients. For example, Prophet (Long and Rinard, 2016), ACS (Xiong et al., 2017), Elixir (Saha et al., 2017), Hanabi (Xiong and Wang, 2022), and LIANA (Chen et al., 2022) all use different learning models for patch ingredient selection. More recently, state-of-the-art deep learning techniques have been employed in APR techniques, such as SEQUENCER (Chen et al., 2019), CoCoNut (Lutellier et al., 2020), CURE (Jiang et al., 2021), DLFix (Li et al., 2020), Recoder (Zhu et al., 2021), RewardRepair (Ye et al., 2022a), SelfAPR (Ye et al., 2022b), and many others (Fan et al., 2023; Xia et al., 2023). These approaches suffer from interpretability issues and mostly can repair simple bugs (e.g., single-line bugs). In contrast, our approach generates patches by reusing existing code, which complements them based on our experimental results.
8. Conclusion
@@
Redundancy-based APR techniques have been widely studied and many approaches have been proposed. However, they often generate a large number of incorrect patches while only effectively repairing a small number of the bugs. To enhance the performance of redundancy-based APRs, this paper introduces a novel repair technique called Repatt, which employs a two-level pattern mining process to facilitate more precise patch generation through fine-grained code reuse. Evaluations conducted on Defects4J v1.2 and v2.0 benchmarks demonstrate that Repatt complements existing approaches by effectively repairing many unique bugs and achieving higher patch precision in real-world repair scenarios. Furthermore, we present the first attempt at combining different approaches to improve the SoTA APRs. The promising results obtained from this study encourage further research in this direction.

### v3

@@ VI Discussion
Internal threats to validity. Similar to existing works, our study also face an internal threat of manually verifying plausible patches to identify correct patches. To address this, the first three authors independently performed a careful analysis of each patch and reached a consensus. In particular, we have also made all generated patches publicly available for further inspection and validation.
External threats to validity. The primary external threat to validity lies in the subjects used in our evaluation, and the performance of Repatt may not be generalized to other datasets. To address this, we evaluate Repatt on two widely used datasets – Defects4J v1.2 and v2.0, which together contain over 800 real-world bugs from 17 diverse projects. Our results demonstrate that Repatt is effective and achieves promising performance. In the future, we plan to evaluate Repatt on more datasets to further address this threat. Moreover, our study also share a common external threat with many existing studies – due to the complexity and resource constraints, we reused the results of baselines reported in their papers without reproducing them in our own environment. To address this, we carefully examined the open-source repositories of the baselines and adhere to the common 5-hour time budget setting used in APR. In the future, we plan to conduct a more comprehensive evaluation of both our approach and the baselines a unified environment to further address this threat.
TABLE IV: Comparison of Repatt with Existing Redundancy-Based APR Approaches
Approach Reuse Granularity Patch Generation Technique Context Awareness
GenProg Statement Genetic Programming (GP)-based random search No
AE Statement Systematic search pruned by semantic equivalence No
RSRepair Statement Random application of mutation operators No
ARJA Statement Multi-objective GP with type-matching No
ssFix ASTNode Syntactic search and code transfer YES
SimFix ASTNode Similarity search and pattern mining YES
CapGen ASTNode Context-aware search and ranking of AST node edits YES
SearchRepair Statement Semantic search via SMT constraint solving YES
HERCULES ASTNode Simultaneous repair of abstract sibling locations YES
TransplantFix ASTNode Graph differencing of CFGs to transplant methods YES
Repatt Token Pattern matching using mined token/S-TAC patterns YES
VII Related Work
@@
There are many automated program approaches have been proposed and achieved good results [6, 60, 61, 62, 11, 51, 10, 63, 8, 2, 3, 12, 54, 20, 21], among which redundancy-based APR approaches have been attracted much attention, such as GenProg [6, 17], AE [61, 64], RSRepair [16], ARJA [64], ssFix [9], SimFix [8], CapGen [1], SearchRpepair [65], HERCULES [66] and so on. All these approaches are based on the “plastic surgery hypothesis” [29]. Specifically, Table IV summarizes the major differences among them , including code reuse granularity, patch generation strategies and whether they are context-aware. For example, GenProg reuses code statement using a genetic programming-based random search and is not context aware, ssFix retrieves syntactically similar code of AST and will rename variables to fit current context. SimFix further incorporates history patches to refine the patch space and consider the context. Similarly, CapGen considers the frequency of tokens for patch generation while HERCULES leverages the co-evolution of program elements. Our approach enhances redundancy-based APR by enabling finer-grained code reuse, differentiating it from existing methods. Unlike SimFix and ssFix, which rely on AST differences, our method operates at the token level, capturing more granular code patterns. Compared to CapGen, which considers token frequency without their contextual relationships, our approach mines structured code patterns for patch generation. Additionally, unlike SearchRepair, we avoid costly semantic code searches and uniquely support nonconsecutive token usage patterns, which have been shown to be effective but are not utilized by existing methods. Other APR techniques, such as PAR [14] and TBar [4], improve patch quality through predefined repair templates, while constraint-solving-based techniques [10, 51, 67, 63, 68, 3, 69, 70] rely on symbolic execution and constraint solving, often facing scalability challenges. Different from these approaches, this work aims to improve the performance of redundancy-based APR techniques and does not face the scalability issue.
@@
With the rapid development of deep learning techniques, the latest APR approaches also leverage such techniques for patch generation. Early studies use statistical machine learning algorithms to sort or pick patch ingredients. For example, Prophet [7], ACS [2], Elixir [15], Hanabi [71], and LIANA [72] all use different learning models for patch ingredient selection. More recently, state-of-the-art deep learning techniques have been employed in APR techniques, such as SEQUENCER [44], CoCoNut [73], CURE [21], DLFix [45], Recoder [18], RewardRepair [74], SelfAPR [19], and many others [75, 43]. These approaches suffer from interpretability issues and mostly can repair simple bugs (e.g., single-line bugs). In contrast, our approach generates patches by reusing existing code, which complements them based on our experimental results. The latest LLM-based APR methods further improved repair performance [22, 76, 43, 20, 77]. For example, Repilot [23] integrates CodeT5 [78] with a completion engine to enhance repair performance, while GAMMA [24] utilizes CodeBERT and UniXcoder [79] to fill predefined repair templates. Unlike Repatt, these methods treat LLMs as black boxes and are still facing challenges related to interpretability [80], data leakage [81], and generalizability [82]. In addition, according to our evaluation results, our approach also complements these LLM-based methods, and thus can be further combined with them for better APR.
VIII Conclusion
@@
To enhance redundancy-based APR, this paper proposes Repatt, a novel repair technique using a two-level pattern mining process for precise patch generation via fine-grained code reuse. Evaluations on Defects4J v1.2 and v2.0 show that Repatt complements existing methods by repairing unique bugs and improving patch precision in real-world repair scenarios. Furthermore, we present the first attempt at combining different approaches to improve the SOTA APRs. Our promising results encourage further research, and we have open-sourced our data and implementations for replication and exploration [40].
Acknowledgment
We thank the editors and anonymous reviewers for their constructive suggestions to help improve the quality of this paper. This work was supported by the National Key Research and Development Program of China (Grant No. 2024YFB4506300), and the National Natural Science Foundation of China (Grant Nos. 62202324, No. 62322208 and No. 62202040).